I am very bad at understanding dimensions. Will someone help shed some light?
"The length of a rectangle is 2 cm more than twice it's width. If the perimeter of the rectangle is 52 cm, find the dimensions of the rectangle."
Thanks!

Answer by funmath(2933)

If I was with you I'd draw a rectangle. It has four sides in which the opposite sides are of equal length. There are 2 lengths and 2 widths. When we find perimeters we simply add all the sides.
Perimeter=length+length+width+width
Let perimeter=P and length =L and width=W and we can find the formula for the perimeter of a rectangel:
P=L+L+W+W
P=2L+2W
In this case we were told that p=52.
We weren't told anything about the width so let W=x.
We were told that the length (L) is (=) 2 more than (+2) twice the width (2x). So let L=2x+2
Substitute the values into the formula P=2L+2W and you have:
52=2(2x+2)+2(x)
52=2(2x)+2(2)+2x
52=4x+4+2x
52=(4+2)x+4
52=6x+4
52-4=6x+4-4
48=6x
48/6=6x/6
8=x
So the width :x=8 cm
The length:2x+2=2(8)+2=16+2=18 cm